The problem is:
Someone paddled upstream at 2mi/h relative to the riverbank. Returning downstream with the current, the speed was 3mi/h.
Find the paddling speed in still water, and the speed of the rivers current.
Could you please help me with this one?
Thank you

Answer by rvquartz(19)

let x = speed of river current
let y = your paddling speed

when going upstream (fighting against the river current) your ACTUAL speed is
(your paddling speed) MINUS (speed of river current)

when going downstream (fighting with the river current) your ACTUAL speed is
(your paddling speed) PLUS (speed of river current)

so,

upstream:
y - x = 2


downstream:
y + x = 3

from upstream equation, we can solve for y and get y = (2 + x)
we can substitute this into the downstream equation and we have
(2 + x) + x = 3

solving for x:
2 + 2x = 3
2x = 1
x = 0.5, and since y = (2 + x), y = 2.5


speed of river is 0.5 miles per hour
your paddling speed (in still water) is 2.5 miles per hour